A bond was purchased on March 20, 2009 and the quoted bond price was $1075. The previous coupon date was January 1, 2009. The next coupon date is January 1, 2010. The bond will mature on January 1, 2018. The bond’s annual coupon rate is 10% and the face value of the bond is $1,000. Coupons will be paid annually.

a. Compute the bond’s yield to maturity on an accrued interest payment basis.

PLEASE USE EXCEL
Answer #1

Interest accrued for period between Jan 1, 2009 to Mar 20, 2009 = 1000 *10% * 79/ 365 = 21.644

Present value of the bond = 1075+ 21.644= 1096.644

Using excel function rate as below to compute YTM:

NPER 9
FV 1000
PMT 100.00 [1000*10%]
PV 1096.644
Rate 8.43% [Rate ( nper, pmt,-pv,fv)]

Any doubt please comment